OpenStudy (anonymous):

Suppose A and B are dependent events. If P(A|B)=0.25 and P(B)=0.8 , what is P(AuB) ?

hartnn (hartnn):

Hi Panda, do you know the formula for P(A|B) ?

OpenStudy (anonymous):

P(AuB)/P(B) ?

hartnn (hartnn):

close, its actually \(P (A|B) = \dfrac{P(A \cap B)}{P(B)}\)

OpenStudy (anonymous):

if i did it right wouldn't it be 0.25, and yeah i know i just didn't know how to set it up on the computer

hartnn (hartnn):

from that \(P (A \cap B) = 0.25 \times 0.8 =..\) isn't it ?

hartnn (hartnn):

also, any other information given ?? like P(A) ?

OpenStudy (anonymous):

well the answers are 0.25, 0.85, 0.8, and 0.2 I'm not quite sure ow to get the right answer :/

OpenStudy (anonymous):

and no theres not thats it

hartnn (hartnn):

we now have \(P(A\cap B)\) using the formula \(\Large P (A\cup B) = P(A)+P(B)-P(A\cap B)\) we still need P(A)

OpenStudy (anonymous):

I'm trying to figure it out and i don't understand that's all the info i have isn't P(A)=0.25

hartnn (hartnn):

if A,B are *independent*, then we can solve it...

hartnn (hartnn):

how is P(A) =0.25 ?

OpenStudy (anonymous):

i don't knowI'm just trying to figure things out and the answer was 0.2

hartnn (hartnn):

maybe the question asked is \(P(A \cap B)\) which we already got as \(P (A \cap B) = 0.25 \times 0.8 =0.2\)

OpenStudy (anonymous):

wow.. see i got that but i scewed up by doing more than i needed to :/ thanks for the help tho.

hartnn (hartnn):

welcome ^_^